OpenStudy (anonymous):

ln(4x-1)=36 solve for x ln(4x-1)=36 solve for x @Mathematics

OpenStudy (anonymous):

Take the e of both sides: 4x-1=e^36 add the 1 and divide by the 4: x=(e^36-1)/4

OpenStudy (anonymous):

Sorry that should be x=(e^36+1)/4

OpenStudy (anonymous):

awesome